# Progressive Data Dropout: An Embarrassingly Simple Approach to Faster Training

### v2

@@ 1 Introduction
Advances in deep learning over the last years have consistently come at the cost of substantial computational resources. Each year, the computational cost of the most notable models multiplies by over 4 times  [^1] https://epoch.ai/trends. This poses challenges both in terms of real-world applications and sustainability Gowda et al. (2023).
@@
Beyond the cognitive inspiration, our approach offers significant practical benefits. By initially focusing on difficult examples, the model processes substantially fewer data points during early epochs, leading to reduced computational requirements. This selective focus not only accelerates training but also appears to guide the model toward more generalizable representations, as evidenced by improved performance at test time. Our experimental results demonstrate that models trained with our approach achieve comparable or superior accuracy while requiring only 0.124$\times$ the number of effective epochs compared to standard training procedures. This efficiency gain is particularly notable for already-efficient architectures like MobileNet and EfficientFormer, where further reducing training costs can significantly impact deployment scenarios with constrained resources. We plot some of these results on CIFAR-100 Krizhevsky and Hinton (2009) in Figure 2.
@@ 3.5 Effective Epochs
Effective Epochs=Number of Samples through Backpropogation over entire runTotal Number of Training $\text{Effective Epochs}=\frac{\text{Number of Samples through Backpropogation over entire run}}{\text{Total Number of Training Samples}}$
@@ 5.1 Datasets
We conduct all classification experiments using three standard image classification benchmarks: CIFAR-10 Krizhevsky and Hinton (2009), CIFAR-100 Krizhevsky and Hinton (2009), and ImageNet Deng et al. (2009). The CIFAR-10 and CIFAR-100 datasets each consist of $60{,}000$ $32\times 32$ color images, split into $50{,}000$ training and $10{,}000$ test images. CIFAR-10 contains 10 object classes, while CIFAR-100 spans a more fine-grained taxonomy of 100 classes. Both datasets are widely used to benchmark small and mid-sized neural architectures. We also use the ImageNet (ILSVRC2012) dataset, which contains over $1.28$ million training images and $50{,}000$ validation images across $1{,}000$ object categories. The images are of higher resolution and greater diversity compared to CIFAR datasets, making ImageNet a robust benchmark for evaluating both model capacity and training efficiency. In addition to supervised classification, we also evaluate our method in a self-supervised learning setup, using ImageNet as the benchmark dataset. This allows us to assess the generality of Progressive Data Dropout beyond supervised learning scenarios.
@@ 5.3 Supervised Image Classification
EfficientNet-B0 88.21 / 30 87.82 / 5.86 90.45 / 15.71 87.80 / 6.70 -0.41 4.48 $\times$
MobileNet 85.49 / 30 85.39 / 6.42 88.24 / 15.71 86.12 / 7 0.63 4.29 $\times$
ResNet-50 84.85 / 30 85.61 / 5.08 87.34 / 15.71 86.10 / 6.98 1.25 4.30 $\times$
EfficientFormer-L1 80.32 / 30 81.17 / 5.82 81.12 / 15.71 81.95 / 10.74 1.63 2.79 $\times$
@@
EfficientNet-B0 77.10 / 350 77.45 / 111.22 64.71 / 20 79.51 / 142.44 2.41 2.46 $\times$
MobileNet-V2 71.60 / 250 68.42 / 54.58 63.85 / 20 73.25 / 98.03 1.75 2.55 $\times$
ResNet-50 75.04 / 100 76.21 / 33.02 71.54 / 20 79.18 / 46.44 4.14 2.15 $\times$
EfficientFormer-L1 79.11 / 300 79.24 / 91.48 69.67 / 20 80.25 / 106.07 1.14 2.83 $\times$
ViT-B-MAE 83.10 / 100 83.31 / 35.36 79.51 / 20 84.13 / 41.75 1.03 2.40 $\times$
@@ 5.4 Self-Supervised Training
Despite its simplicity, SRD reduces the total number of effective epochs from 800 to just 50, a 16$\times$ reduction in training steps. This reduction is especially impactful given the extreme time, memory, and energy requirements of MAE pretraining. Moreover, across all downstream tasks, our method remains highly competitive with the baseline, with only negligible loss in performance. We report these results in Table 3.
@@
EE Accuracy Accuracy APbox
@@
SRD 0.98 49.95 82.65% 86.18% 49.6
Table 3: Baseline trained to match the effective number of epochs (EE) used by dropout variants.
Variant EE Baseline (%) Accuracy (%)
@@ 5.5 Ablation Study
How sensitive is difficulty-based dropout to the choice of threshold? We vary the confidence threshold $\tau$ used for dropping samples in the difficulty-based variant (DBPD), ranging from 0.1 to 0.9 in increments of 0.1 and report these results in Table 4. Higher thresholds lead to more conservative early-stage dropout, while lower values remove more data per epoch. Hence, lower thresholds tend to have a much lower step count. To ensure a balance between accuracy and epochs, we use thresholds of 0.3 and 0.7 in all experiments. All experiments are conducted using EfficientNet-B0 on CIFAR-100 for 200 epochs with the final epoch being full-dataset revision.
@@
How many samples are we dropping over time? Figure 4 illustrates the number of training samples retained at each epoch under the DBPD (Difficulty-Based Progressive Dropout) strategy for two different threshold settings: 0.3 (darker line) and 0.7 (lighter line). In both cases, training begins with the full dataset. As the model improves and classifies more samples with high confidence, the number of retained examples drops sharply, particularly in the early epochs. The higher threshold (0.7) causes a more aggressive pruning of data, dropping a larger fraction of confident samples per epoch. As a result, it reaches a low sample count much earlier, within the first 15–20 epochs, leaving fewer than 10,000 samples for much of training. In contrast, the 0.3 threshold maintains more samples for a longer duration, leading to a more gradual decay. Both strategies conclude with a final epoch in which the full dataset is reintroduced for revision, producing a sharp spike back to 50,000 samples. This U-shaped trajectory reflects the design principle of DBPD: prioritize hard examples early, reduce redundant computation mid-training, and consolidate knowledge through a final full-dataset pass. The comparison highlights the trade-off between aggressive data reduction and sustained exposure to diverse samples.
@@ 6 Limitations
While Progressive Data Dropout demonstrates significant improvements in training efficiency and generalization, it has certain limitations. The method is most effective when training models from scratch; in fine-tuning or transfer learning scenarios, the relative gains are slightly lower. Whilst there are still some gains, pre-training helps to reduce that. Our current evaluation is limited to classification tasks, with preliminary extensions of a pre-training approach to object detection. However, broader applicability to other domains such as segmentation, language modeling, or multi-modal learning remains to be investigated. Additionally, we observe that larger datasets generally require more effective epochs and show lower gains, which may limit some of the training savings. Finally, the method introduces hyperparameters such as thresholds or decay rates, that can be sensitive to the dataset and model. While our default settings perform well across multiple benchmarks, optimal tuning may be necessary in more specialized contexts. Future work would be coming up with mathematical approximations to avoid the use of difficulty based scheduling and adapting the approach to different modalities and tasks.
@@ Appendix B Algorithm
Input: Model $f_{\theta}$ with parameters $\theta$, Training data $\mathcal{D}_{\text{train}}$, test data $\mathcal{D}_{\text{test}}$, Number of epochs $E$, revision start epoch $E_{\text{rev}}$, Confidence threshold $\tau:0\leq\tau\leq 1$, Loss function $\mathcal{L}$Decay factor : $10<\gamma<$ 1
for epoch $=1$ to $E$ do
@@ Appendix C Mathematical Approximation for SMRD
Here, $T$ is the total number of training epochs (e.g., $T=200$), and $\hat{f}(x;\alpha)$ is one of the following parameterized decay functions, each normalized to output a maximum of $50{,}000$ (Using CIFAR-100 as our example) samples:
@@
Appendix D Evaluating the approximations on CIFAR-100

### v3

# Progressive Data Dropout: An Embarrassingly Simple Approach to Train Faster

@@ 1 Introduction
footnotetext: * Equal Contribution.
Advances in deep learning over the last years have consistently come at the cost of substantial computational resources. Each year, the computational cost of the most notable models multiplies by over 4 times Epoch AI (2023). This poses challenges both in terms of real-world applications and sustainability Gowda et al. (2023).
@@
[…]
@@ 3.5 Effective Epochs
[…]
@@ 5.1 Datasets
We conduct all classification experiments using three standard image classification benchmarks: CIFAR-10 Krizhevsky and Hinton (2009), CIFAR-100 Krizhevsky and Hinton (2009), and ImageNet Deng et al. (2009). The CIFAR-10 and CIFAR-100 datasets each consist of 60,00060{,}000 $32\times 32$ color images, split into 50,00050{,}000 training and 10,00010{,}000 test images. CIFAR-10 contains 10 object classes, while CIFAR-100 spans a more fine-grained taxonomy of 100 classes. Both datasets are widely used to benchmark small and mid-sized neural architectures. We also use the ImageNet (ILSVRC2012) dataset, which contains over $1.28$ million training images and 50,00050{,}000 validation images across 1,0001{,}000 object categories. The images are of higher resolution and greater diversity compared to CIFAR datasets, making ImageNet a robust benchmark for evaluating both model capacity and training efficiency. In addition to supervised classification, we also evaluate our method in a self-supervised learning setup, using ImageNet as the benchmark dataset. This allows us to assess the generality of Progressive Data Dropout beyond supervised learning scenarios.
@@ 5.3 Supervised Image Classification
EfficientNet-B0 90.41 / 200 88.22 / 18 92.07 / 84 90.55 / 20 0.14 10 $\times$
MobileNet-V2 88.46 / 200 89.55 / 22.3 90.73 / 84 90.91 / 24.0 2.45 8.33 $\times$
ResNet-50 87.94 / 200 88.42 / 23 89.81 / 84 89.12 / 22.1 1.18 9.05 $\times$
EfficientFormer-L1 82.38 / 200 84.47 / 28.2 84.38 / 84 85.20 / 31 2.82 6.45 $\times$
@@
EfficientNet-B0 77.10 / 350 77.45 / 111.2 64.71 / 20 79.51 / 142.4 2.41 2.46 $\times$
MobileNet-V2 71.60 / 250 68.42 / 54.5 63.85 / 20 73.25 / 98.0 1.75 2.55 $\times$
ResNet-50 75.04 / 100 76.21 / 33.0 71.54 / 20 79.18 / 46.4 4.14 2.15 $\times$
EfficientFormer-L1 79.11 / 300 79.24 / 91.5 69.67 / 20 80.25 / 106.1 1.14 2.83 $\times$
ViT-B-MAE 83.10 / 100 83.31 / 35.4 79.51 / 20 84.13 / 41.7 1.03 2.40 $\times$
@@ 5.4 Self-Supervised Training
[…]
@@
$E$ Accuracy Accuracy APbox
@@
SRD 0.98 49.95 77.89% 83.81% 47.2
Table 3: Baseline trained to match the effective number of epochs ($E$) used by dropout variants.
Variant $E$ Baseline (%) Accuracy (%)
@@
5.5 Comparison to State-of-the-Art
We conducted comparisons against a suite of existing methods, including DataDiet (ELN and Forget) Paul et al. (2021), IES Yuan et al. (2025), Early Stopping, and InfoBatch Qin et al. . We followed the official implementations for all baselines and used EE to evaluate efficiency. Our method outperforms all compared approaches in both accuracy and EE, with higher values indicating better performance for both metrics. The results, reported as Accuracy / EE, are obtained by fine-tuning models on CIFAR-100 (pretrained on ImageNet) for 200 epochs.
Table 4: Comparison with strong active learning baselines on CIFAR-100.
Model Baseline DataDiet ELN DataDiet Forget IES Early Stopping InfoBatch SMRD (ours)
ResNet-50 78.5 / 200 78.6 / 141 79.1 / 142 79.5 / 157 74.3 / 121 80.4 / 145 81.5 / 33
EfficientNet 83.3 / 200 82.7 / 139 83.0 / 139 83.7 / 153 76.6 / 132 84.0 / 141 84.4 / 34
EfficientFormer-L1 85.3 / 200 84.8 / 142 85.0 / 140 85.8 / 155 78.2 / 125 86.6 / 143 87.5 / 51
5.6 Generalization in Language Tasks
We used GPT-2 pre-training with SRD on the FineWeb-Edu dataset Penedo et al. (2024). The results in the Table 5 show that SRD achieved substantial computational savings, reducing both the number of training steps and wall-clock time by more than half.
Table 5: NLP task performance of SRD for small GPT-2 pretraining experiments.
Model PPL_val PPL_train Num_steps Wall clock
GPT-2 (Baseline) $29.37$ $28.50$ $2,560,000$ $13.34$ hrs
SRD (Threshold = 0.99) $32.81$ $31.78$ $\mathbf{1,066,720}$ $\mathbf{5.91}$ hrs
@@ 5.7 Ablation Study
How sensitive is difficulty-based dropout to the choice of threshold? We vary the confidence threshold $\tau$ used for dropping samples in the difficulty-based variant (DBPD), ranging from 0.1 to 0.9 in increments of 0.1 and report these results in Table 6 , Table 7, and Table 8. Higher thresholds lead to more conservative early-stage dropout, while lower values remove more data per epoch. Hence, lower thresholds tend to have a much lower step count. To ensure a balance between accuracy and epochs, we use thresholds of 0.3 and 0.7 in all experiments. Experiments in Table 6 are conducted using EfficientNet-B0 on CIFAR-100 for 200 epochs with the final epoch being full-dataset revision. For SRD, we applied it in the most straightforward form of data dropping. Table 9 presents the results on CIFAR-100 across different thresholds for reference.
@@
Table 7: Analysis for different dropout thresholds on CIFAR-100 dataset with training from scratch.
Threshold 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
EfficientFormer $55.91$ $56.37$ $56.88$ $57.52$ $57.85$ $57.98$ $58.12$ $58.08$ $58.17$ $58.22$
MobileNet-v2 $61.25$ $61.49$ $62.17$ $62.85$ $62.98$ $63.21$ $63.27$ $63.31$ $63.86$ $63.79$
Table 8: Analysis for different dropout thresholds on ImageNet with training from scratch.
Threshold 0.0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9
ResNet-50 $75.04$ $75.55$ $75.80$ $76.21$ $76.34$ $76.42$ $76.79$ $77.14$ $77.20$ $77.19$
Table 9: SRD Threshold sensitivity on CIFAR-100 dataset.
Threshold 0.0 0.2 0.4 0.6 0.8 0.9 0.92 0.94 0.95 0.96
EfficientFormer $55.91$ $17.84$ $27.92$ $38.85$ $48.83$ $56.15$ $56.74$ $56.91$ $56.95$ $56.97$
MobileNet-v2 $61.25$ $24.41$ $35.22$ $48.63$ $57.79$ $62.18$ $62.59$ $62.77$ $62.80$ $62.82$
@@
How many samples are we dropping over time? Figure 4 illustrates the number of training samples retained at each epoch under the DBPD (Difficulty-Based Progressive Dropout) strategy for two different threshold settings: 0.3 (darker line) and 0.7 (lighter line). In both cases, training begins with the full dataset. As the model improves and classifies more samples with high confidence, the number of retained examples drops sharply, particularly in the early epochs. The higher threshold (0.7) causes a more aggressive pruning of data, dropping a larger fraction of confident samples per epoch. As a result, it reaches a low sample count much earlier, within the first 15–20 epochs, leaving fewer than 10,000 samples for much of training. In contrast, the 0.3 threshold maintains more samples for a longer duration, leading to a more gradual decay. Both strategies conclude with a final epoch in which the full dataset is reintroduced for revision, producing a sharp spike back to 50,000 samples. This U-shaped trajectory reflects the design principle of DBPD: prioritize hard examples early, reduce redundant computation mid-training, and consolidate knowledge through a final full-dataset pass.
@@ 6 Limitations
While Progressive Data Dropout demonstrates significant improvements in training efficiency and generalization, it has certain limitations. The method is most effective when training models from scratch; in fine-tuning or transfer learning scenarios, the relative gains are slightly lower. Whilst there are still some gains, pre-training helps to reduce that. Our current evaluation is limited to classification tasks, with preliminary extensions. However, broader applicability to other domains such as segmentation, language modeling, or multi-modal learning remains to be investigated. Additionally, we observe that larger datasets generally require more effective epochs and show lower gains, which may limit some of the training savings. Finally, the method introduces hyperparameters such as thresholds or decay rates, that can be sensitive to the dataset and model. While our default settings perform well across multiple benchmarks, optimal tuning may be necessary in more specialized contexts. Future work would be coming up with mathematical approximations to avoid the use of difficulty based scheduling and adapting the approach to different modalities and tasks.
@@ Appendix B Algorithm
Input: Model $f_{\theta}$ with parameters $\theta$, Training data $\mathcal{D}_{\text{train}}$, test data $\mathcal{D}_{\text{test}}$, Number of epochs $E$, revision start epoch $E_{\text{rev}}$, Confidence threshold $\tau:0\leq\tau\leq 1$, Loss function $\mathcal{L}$Decay factor : $0<\gamma<1$
for epoch =1=1 to $E$ do
@@ Appendix C Mathematical Approximation for SMRD
Here, $T$ is the total number of training epochs (e.g., $T=200$), and $\hat{f}(x;\alpha)$ is one of the following parameterized decay functions, each normalized to output a maximum of 50,00050{,}000 (Using CIFAR-100 as our example) samples:
@@
Appendix D Evaluating the approximations on CIFAR-10 and CIFAR-100
@@
Do we still have consistent improvements if we are having a longer schedule? We conducted additional experiments using longer training runs (200 epochs) on CIFAR-10 with EfficientNet-B0, ResNet-50, and MobileNet-v2 architectures. To ensure fairness, we used uniform optimization parameters across all models. Table 13 show that our method achieves consistent improvements under the longer training schedule.
Table 13: Fine-tuning models pretrained on ImageNet (200 epochs).
Model Baseline DBPD (0.3) SRD (0.99) SMRD (0.3)
EfficientNet-B0 95.55 / 200 96.11 / 25 96.40 / 84 96.83 / 31
ResNet-50 96.41 / 200 96.73 / 27 96.83 / 84 97.71 / 33
MobileNet-v2 93.91 / 200 94.41 / 26 95.12 / 84 95.55 / 33
What happens if we use strong regularization or augmentation method? Using stronger regularization or augmentation reduces the impact of the current threshold settings in our method. As mentioned in the paper, we report results with a threshold of 0.3 to achieve a balance between speed and accuracy. However, if accuracy is prioritized, higher thresholds can be used, as demonstrated in the table. These results are based on CIFAR-10 training for 200 epochs, using CutMix style augmentation with the parameter $\alpha$ set to 1.
Table 14: Performance under CutMix augmentation.
Model Baseline DBPD(0.3) DBPD(0.8) SRD(0.99) SMRD(0.3) SMRD(0.8)
MobileNet v2 $90.41/200$ $90.96/22$ $91.45/32$ $\mathbf{93.29/84}$ $91.41/26$ $93.24/42$
How does increasing revision epochs or extending total epochs affect? As shown in Figure 4, adding more revision epochs causes overfitting and can degrade test accuracy. To validate this further, we tested SMRD with 25, 50, and 75 revision epochs (compared to the standard setting of 1). While these settings brought the accuracy closer to the baseline, they did not yield any improvements and risked overfitting as shown in Table 16. Additionally, we increased the overall number of training epochs to bring the overall computation closer to the baseline. As shown in Table 16, this did not really improve our performance much.
Table 15: Results of SMRD by increasing revision epochs.
Method Revision Epochs Accuracy (%)
SMRD (standard) 11 $71.01$
SMRD $25$ $68.55$
SMRD $50$ $66.35$
SMRD $75$ $66.27$
Baseline $66.32$
Table 16: Results of extending total training epochs.
Method Total Epochs Accuracy (%) Effective Epochs (EE)
DBPD $200$ $\mathbf{67.15}$ $25$
DBPD $300$ $67.09$ $29$
DBPD $400$ $66.28$ $33$
DBPD $500$ $66.59$ $36$
SMRD $200$ $\mathbf{71.01}$ $35$
SMRD $300$ $70.82$ $44$
SMRD $400$ $70.75$ $49$
SMRD $500$ $70.99$ $51$
Baseline $200$ $66.32$ $200$
Baseline $300$ $\mathbf{66.45}$ $200$
Baseline $400$ $66.37$ $200$
Baseline $500$ $66.35$ $200$
What happens if the dataset has class imbalance or noisy labels? It is important to emphasize that PDD fundamentally focuses on computational efficiency rather than robustness to class imbalance or noise. Table 18 presents results for Long-Tail Classification on CIFAR-10 with $\rho$ = 100, where $\rho$ denotes the ratio between the sample sizes of the most frequent and least frequent classes. The dataset is generated following the same procedure as LDAM-DRW Cao et al. [2019]. We use Focal Loss Lin et al. [2017], consistent with the baseline, to ensure a fair comparison. From the table, all PDD variants show substantial gains over the baseline, with SMRD (0.3) emerging as the best-performing method under class imbalance, achieving the highest accuracy. However, compared to the standard setting, this configuration requires more EE. We evaluated our method on noisy label classification using the Noisy CIFAR-10 dataset, following the setup from Wei et al. [2021], with a noise rate of 17.3% and random noise type, training for 50 epochs. Table 18 shows that all three PDD variants outperform the baseline in both accuracy and efficiency. Notably, SRD (0.95) achieves the highest accuracy, though with slightly higher EE. SMRD and DBPD provide a better trade-off, offering much lower EE while maintaining strong performance, demonstrating their robustness under label noise.
Table 17: Long-tail classification of ResNet-34 on CIFAR-10.
Model Accuracy (%) Effective Epochs (EE)
Baseline 46.04 200
DBPD (0.3) 58.09 64
DBPD (0.2) 54.82 55
DBPD (0.1) 24.28 65
SRD (0.95) 54.28 84
SMRD (0.3) 60.14 58
Table 18: Results of noisy label classification on Noisy CIFAR-10
Model Accuracy (%) Effective Epochs (EE)
Baseline 71.91 50
DBPD (0.3) 74.26 9
SRD (0.95) 75.23 18
SMRD (0.3) 73.76 9
What is the ideal dataset characteristics for PDD? PDD is particularly beneficial for mid-sized to large datasets, where the computational savings are most significant. We also evaluated it on smaller fine-grained datasets, such as CUB and Flowers, and observed consistent computational savings without any loss in accuracy as shown in Table 19. These results highlight PDD’s suitability across different dataset scales and complexities.
Table 19: Model performance comparison on smaller fine-grained datasets.
Model Dataset SMRD (0.3) Baseline
EfficientNet-B0 Airplane 80.07 / 9 78.06 / 100
EfficientNet-B0 CUB 71.03 / 9 69.83 / 100
MobileNet-v2 Airplane 76.92 / 10 73.87 / 100
MobileNet-v2 CUB 68.94 / 10 67.45 / 100
How do the three variants perform given the same fixed compute budget? We conducted a preliminary experiment on CIFAR-100 using ResNet-50, aligning the EE across all methods. All models were fine-tuned from ImageNet pre-training. Table 20 shows that even under equal EE, all PDD variants outperform the baseline, with SMRD remaining the best-performing variant.
Table 20: Performance of aligning Effective Epochs (EE) across methods on CIFAR-100 (ResNet-50)
Model DBPD SRD SMRD
ResNet-50 79.92 / 21 / 0.3 79.11 / 20.8 / 0.93 80.61 / 20.3 / 0.23
What is the practical usability of SMRD? To demonstrate the practical usability of SMRD, we derived simple mathematical approximations of the dropout schedule that closely replicate the oracle SMRD performance (Appendix Section C). Among these, the logarithmic decay approximation provides the closest match. Empirical results in Table 21 show that these approximations still clearly outperform the baselines, confirming their practicality. Moreover, SRD and DBPD also outperform the baselines in most scenarios, further validating their effectiveness and applicability.
Table 21: Practical Usability of SMRD with Logarithmic Decay Approximation.
Model Variant Accuracy Effective Epochs
ResNet-50 SMRD (oracle) 81.54 33
ResNet-50 SMRD (log decay) 81.11 37
EfficientNet-B0 SMRD (oracle) 84.45 34
EfficientNet-B0 SMRD (log decay) 83.92 36
MobileNet-v2 SMRD (oracle) 74.73 36
MobileNet-v2 SMRD (log decay) 74.36 38
What is the wall-clock time using PDD? While our primary metric, Effective Epochs (EE), enables hardware-independent comparison, we also provide several wall-clock time savings estimates in Table 22 for reference. These experiments were conducted on a single NVIDIA RTX 4060 GPU with 8 GB of RAM.
Table 22: Wall clock time savings estimates.
Model Baseline DBPD (0.3) SRD (0.95) SMRD (0.3)
EfficientNet-B0 (CIFAR-100) 15.74 hr 7.89 hr 6.98 hr 9.95 hr
MobileNet-v2 (CIFAR-100) 12.62 hr 7.11 hr 6.67 hr 9.51 hr